OpenStudy (anonymous):

2/5 (12+x) = 6- 1/4 (9-x) Must solve

OpenStudy (kl0723):

\[\frac{ 2 }{ 5 }(12+x) = 6-\frac{ 1 }{ 4 }(9-x)\]

OpenStudy (kl0723):

like this?

OpenStudy (anonymous):

Yes thats right :) I must solve

OpenStudy (anonymous):

u have to multiply the WHOLE equation with 5 and 4. so that, ur equation will not have any fraction number anymore. do u get my point? if u already done it. then, what is ur answer?

OpenStudy (kl0723):

that's what I get :)

OpenStudy (anonymous):

Perfect!!! I kept trying to work with the fractions and it seemed fruitless. Thank you so much!!! Appreciate all the help! @kl0723

OpenStudy (kl0723):

yw :)

OpenStudy (anonymous):

x = -7 Refer to the Mathematica attachment.

OpenStudy (anonymous):

@robtobey Hello! How do you achieve the -7?

OpenStudy (anonymous):

Multiply each side of the equation by 20, "4 times 5".\[\frac{20}{5} (2 (x+12))=20 \left(6-\frac{9-x}{4}\right) \]\[8 x+96=5 x+75 \]Can you complete the solution from here?

OpenStudy (anonymous):

Some of the terms may not appear exactly as those written in the original problem statement, but they are still valid.

OpenStudy (anonymous):

\[3 x=75-96 \]\[3 x=-21 \] x = -7

OpenStudy (anonymous):

I am still so confused. How are you getting 75? @robtobey

OpenStudy (anonymous):

Hold on. I will post a step y step solution.

OpenStudy (anonymous):

Thank you so much! @robtobey

OpenStudy (anonymous):

A detailed solution using Mathematica 9 Home Edition is attached. Sorry it took so long to put it together.

OpenStudy (anonymous):

@robtobey I finally understand!!!! You are great!

OpenStudy (anonymous):

You are welcome.